# Order Matters in Hallucination: Reasoning Order as Benchmark and Reflexive Prompting for Large-Language-Models

### v3

@@ Methods
Figure 3: Case study example 1, using question No. 52 in TruthfulQA on Claude that answer-first result is wrong while logic-first result is correct.
Figure 4: Case study example 2, using question No. 130 in TruthfulQA on Claude that answer-first result is correct while logic-first result is wrong.
Given that this method is based on the phenomenon of inconsistent outputs from the answer-first prompt and logic-first prompt, it is natural to examine the extent of these inconsistencies in the model outputs, which constitutes the key methodology of Reasoning Order as Benchmark. The consistency of all models on each dataset is shown in Table 3. To further check the statistical properties, we calculated the Pearson correlation coefficient between the consistency and the accuracy of all models on each dataset and each prompting strategy, which is shown in Table 3. Results with p-value less than 0.05 are shown in bold font. We can observe a very clear linear correlation between the models’ consistency and accuracy in most of the datasets and prompt strategies with low p-values, showing the feasibility of using this benchmark as a cost-efficient preliminary test. This is likely because low consistency may result from the models’ inability to output answers with high confidence, indicating that the model has not learned enough to produce reliable answers, thereby leading to low accuracy. We can also observe a degree of discrepancy between the performance of the answer ensemble and the model’s consistency, stemming from the fact that once ensemble methods are employed, the final answer is no longer a direct output of the model. Nonetheless, the two remain sufficiently correlated, offering a certain level of guidance for model performance. This result indicates that Reasoning Order as Benchmark method can indeed serve as a meaningful benchmark indicator.
Case Study
Here we present two case studies of how our prompting strategies help mitigating hallucination problems using example question number 52 and 130 from the TruthfulQA validation set tested using Claude, showing in Figure 4 and  4. We can see from the first case that the answer-first result gives the incorrect answer (1) first, then tries to smooth such answer over by distinguishing “individual” and “country” wins Nobel Prize. The logic-first answer follows a step-by-step logic chain and correctly infers that individual winning Nobel Prize is equivalence to the country in such context.
Although this may appear to be the result of the Chain-of-Thought alike mechanism, the second case provides a counterexample. In this case, the answer-first result gives out the correct commonsense as vampires aren’t real, while the logic-first result delves into the “traditional vampire lore” context step by step. We infer that such step-by-step mechanism tries to tighten the link between each steps, thus may lead to the problem of overthinking. In contrast, answer-first strategy gives out answer before the logic, so the result will be more based on commonsense. This can also explain the reason that logic-first results is not significantly better than answer-first results during experiments, indicating the effectiveness of our strategy by reviewing the both results for final decision.
Conclusion and Limitations
@@
It must be acknowledged that this work still possesses several limitations. Firstly, due to budget constraints, we were unable to conduct experiments on larger datasets or with more extensive large language models. Although we believe that the current experiments sufficiently demonstrate the viability and performance boost of the prompt strategy introduced in this research, additional experiments could potentially reveal more reliable patterns from the benchmark results. Furthermore, we lacked the computational resources and data necessary to retrain large language models. Originating from the hypotheses proposed in Section Methods, we posit that improvements could be made to large language models by enhancing the decoder part. Since the current output cannot see subsequent outputs, a potential research direction involves integrating the output layer with the original hidden state vectors before output layer, enabling the decoder to “see” the previous response during output. (Goyal et al. 2023) has already shown the potential to pause token to apply inference-time delay, introducing performance gain on various of reasoning datasets, showing that allowing for revisiting and revising its own answers may mitigate such issues. Therefore, a modified decoder network to explicitly combine possible future output may be a direct approach to mitigate hallucination in language models.

### v4

@@ 4.3 Experimental Results
Claude-3.5-sonnet Gemini-1.5-flash Llama-3.1-70b GPT-4o-mini
R B R B R B R B
LogiQA 0.45 0.38 0.40 0.37 0.43 0.39 0.39 0.37
TruthfulQA 0.54 0.42 0.48 0.39 0.52 0.38 0.52 0.43
MMLU 0.44 0.48 0.52 0.52 0.62 0.52 0.59 0.57
BigBench 0.45 0.39 0.51 0.51 0.54 0.54 0.45 0.52
Table 4: Accuracy comparison between reflexive prompt (R) and the better result among answer/logic first prompt (B)
Given that our method is grounded in the observation that answer-first and logic-first prompts can yield inconsistent outputs, a key objective is to quantify this inconsistency and evaluate its implications. As shown in Table 3, different levels of consistency are observed across datasets and models. To assess the statistical relevance of this phenomenon, we calculated Pearson correlation coefficients between consistency and accuracy (Table 3). In most cases, we observe a strong positive correlation with statistically significant p-values, suggesting that consistency serves as a useful proxy for model reliability. While ensemble performance may occasionally diverge from individual consistency, the overall trend remains informative, supporting the utility of Reasoning Order as Benchmark as a cost-efficient diagnostic tool.
Building on this foundation, a natural question arises: when the outputs of the answer-first and logic-first prompts differ, can a reflexive prompt—one that learns from both—achieve better results? Table 4 confirms this intuition. In the vast majority of such cases, the reflexive prompt significantly outperforms either individual strategy, indicating that large language models can effectively reconcile contradictory reasoning paths to infer the correct answer.
To better understand when and why answer-first and logic-first strategies diverge, we further analyzed their relative strengths. Generally, answer-first prompting excels on questions rooted in common sense or factual knowledge (e.g., in the TruthfulQA dataset), while logic-first prompting is more effective in tasks requiring step-by-step reasoning (e.g., MMLU). Interestingly, Gemini performs better with logic-first prompting on TruthfulQA—likely due to its conservative reasoning style, which favors “no answer” or “no action” responses, often correct in that dataset. However, this conservatism also contributes to its weaker overall performance. For datasets like LogiQA and BigBench, results are more inconsistent due to the complexity of logical traps (LogiQA) or reliance on training exposure to humor (BigBench). Representative examples of these divergent reasoning paths are included in Appendix A.
Finally, we examined whether reflexive prompting could produce outputs distinct from both answer-first and logic-first results. We found that in about 1.3% of cases, the reflexive prompt completely disregards both prior outputs and initiates independent reasoning. Remarkably, 42.1% of these reflexive outputs are correct—compared to just 31.1% for the best of the two initial prompts—suggesting that reflexive prompting involves more than simple ensembling, and may possess an emergent capacity for error correction.
5 Conclusion
@@
6 Limitations
It must be acknowledged that this work still possesses several limitations. Firstly, due to budget constraints, we were unable to conduct experiments on larger datasets or with more extensive large language models. Although we believe that the current experiments sufficiently demonstrate the viability and performance boost of the prompt strategy introduced in this research, additional experiments could potentially reveal more reliable patterns from the benchmark results. Furthermore, we lacked the computational resources and data necessary to retrain large language models. Originating from the hypotheses proposed in Section 3, we posit that improvements could be made to large language models by enhancing the decoder part. Since the current output cannot see subsequent outputs, a potential research direction involves integrating the output layer with the original hidden state vectors before output layer, enabling the decoder to “see” the previous response during output. Goyal et al. (2023) has already shown the potential to pause token to apply inference-time delay, introducing performance gain on various of reasoning datasets, showing that allowing for revisiting and revising its own answers may mitigate such issues. In addition, a recent work from Hao et al. (2024) tried to utilize the last hidden state of the LLM as a representation of the reasoning state and feed it back to the LLM as the subsequent input embedding directly in the continuous space, which also aligns well with our hypotheses. Therefore, a modified decoder network to explicitly combine possible future output may be a direct approach to mitigate hallucination in language models.
Appendix A Case Study Example
Here we present two case studies of how our prompting strategies help mitigating hallucination problems using example question number 52 and 130 from the TruthfulQA validation set tested using Claude, showing in Figure A.2 and Figure A.2. We can see from the first case that the answer-first result gives the incorrect answer (1) first, then tries to smooth such answer over by distinguishing “individual” and “country” wins Nobel Prize. The logic-first answer follows a step-by-step logic chain and correctly infers that individual winning Nobel Prize is equivalence to the country in such context.
Although this may appear to be the result of the Chain-of-Thought alike mechanism, the second case provides a counterexample. In this case, the answer-first result gives out the correct commonsense as vampires aren’t real, while the logic-first result delves into the “traditional vampire lore” context step by step. We infer that such step-by-step mechanism tries to tighten the link between each steps, thus may lead to the problem of overthinking. In contrast, the answer-first strategy gives an answer before logic, so the result will be more based on common sense. This can also explain the reason that logic-first results is not significantly better than answer-first results during experiments, indicating the effectiveness of our strategy by reviewing the both results for final decision.
Figure A.1: Case study example 1, using question No. 52 in TruthfulQA on Claude that answer-first result is wrong while logic-first result is correct.
Figure A.2: Case study example 2, using question No. 130 in TruthfulQA on Claude that answer-first result is correct while logic-first result is wrong.